OpenStudy (anonymous):

How do you find the maxima and minima for g(x) = x^2 - 2/x

OpenStudy (aum):

First note that g(x) is not defined when x = 0. So the domain of g(x) excludes x. 0 = 2x + 2/x^2 multiply by x^2 0 = 2x^3 + 2 = 2(x^3 + 1) x^3 = -1 x = -1.

OpenStudy (aum):

I meant - So the domain of g(x) excludes 0.

OpenStudy (aum):

Well, see my note above. x = 0 is excluded from the domain of g(x).

OpenStudy (aum):

But the original function does not exist at x = 0 and so it is excluded from g'(x) or g''(x).

OpenStudy (aum):

The function has to be continuous in an interval. g(x) is not continuous at x = 0. So we exclude x = 0.

terenzreignz (terenzreignz):

You should be... critical numbers are where the *derivative* is either zero or undefined. ^_^

terenzreignz (terenzreignz):

If a function itself is not defined at a point, then there is no...point in pondering whether that undefined point could be a max or a min... am I right? :D

terenzreignz (terenzreignz):

So, I guess that about wraps things up? :D

terenzreignz (terenzreignz):

I should think so ^_^

terenzreignz (terenzreignz):

But the question is about finding extrema... is this critical number an extremum at all? ^^

terenzreignz (terenzreignz):

Fair enough ^_^

terenzreignz (terenzreignz):

A positive second derivative implies upward concavity... just saying ;)

terenzreignz (terenzreignz):

Well, positive second derivative means that it's curving upward, more or less

terenzreignz (terenzreignz):

Precisely ^_^ (this is assuming that the first derivative was zero, mind you ;) )

terenzreignz (terenzreignz):

No problem.

terenzreignz (terenzreignz):

It does look like @aum did most of the heavy lifting though. Just saying... again ;)

OpenStudy (aum):

no big deal. As long as the OP gets what he was seeking for.

terenzreignz (terenzreignz):

The joy of learning... bliss :D

OpenStudy (aum):

I looked up the definition of critical points: We say that x = c is a critical point of the function f(x), IF: f(c) exists AND either f'(c) = 0 OR f'(c) does not exist. http://tutorial.math.lamar.edu/Classes/CalcI/CriticalPoints.aspx

OpenStudy (aum):

Here f(0) does not exist. Therefore, x = 0 is not a critical point.